OpenStudy (anonymous):

What is the theoretical period with a length of a pendulum with 0.477

OpenStudy (anonymous):

Also What is the theoretical period with a length of a pendulum with 0.3785?

OpenStudy (anonymous):

@nubeer @MathLegend @mathawesomeness @mathgeek898

OpenStudy (anonymous):

is it 3???

OpenStudy (anonymous):

.47 what? feet meter?

OpenStudy (anonymous):

seconds

OpenStudy (anonymous):

What is L & G stand for

OpenStudy (anonymous):

L : pendulum length g : acceleration of gravity

OpenStudy (anonymous):

0.477 is NOT in seconds<<<<<<<<

OpenStudy (anonymous):

what is it ??

OpenStudy (anonymous):

minutes maybe, Im not sure.

OpenStudy (anonymous):

is it in [CM]

OpenStudy (anonymous):

no it is Length

OpenStudy (anonymous):

[cm] or [in]

OpenStudy (anonymous):

Its not inches i thought it was in time but it is cm

OpenStudy (anonymous):

i think its in [m] did u convert in the LAB

OpenStudy (anonymous):

no

OpenStudy (anonymous):

im not even sure how to do the last chart im just doing wht i think is right

OpenStudy (anonymous):

how did u come up with that? And did i get the experimental period right?

OpenStudy (anonymous):

is it 0.47 cm ?

OpenStudy (anonymous):

Experimental Period 0.477

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

what is experimatal period for 2?

OpenStudy (anonymous):

@julliemyers Listen to your self, it's likely you say that your height is .....second!!!

OpenStudy (anonymous):

the awsners i had for length pendulum are the same as the experimatal perod?

OpenStudy (agent0smith):

\[T = 2 \pi \sqrt{\frac{ L }{ g }}\] T is period, L is length, g is gravity (9.8m/s^2). This is valid as long as the angle you swing it at isn't too steep.

OpenStudy (agent0smith):

T is in seconds, L in metres.

OpenStudy (anonymous):

the angle was 20 degress

OpenStudy (agent0smith):

20 degrees should be fine for that formula. But what are those lengths 0.477 and 0.3785...? You need units. Below that you have the length as 25.6cm and 12.7cm, so what are 0.477 and 0.3785 supposed to be?

OpenStudy (agent0smith):

20 degrees to the vertical, that is. Like this: 

OpenStudy (agent0smith):

Don't put your times in minutes - use seconds. Things like the average time for 10 swings and period of pendulum should be in seconds.

OpenStudy (agent0smith):

If you're using gravity as 980cm/s^2, then it's okay to use length L in cm.

OpenStudy (anonymous):

Experimental Period 0.477 is a time its not a Length

OpenStudy (agent0smith):

But it's listed in the table under Length of Pendulum...

OpenStudy (anonymous):

but if you chk the work its Time NOT length

OpenStudy (anonymous):

its the sum of the " Period of pendulum "

OpenStudy (agent0smith):

Yeah, that's what I figured, i was about to enter her numbers into the formula. If she's used minutes for the period t, then it's not correct, since gravity is in seconds.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

you can use this site http://tini.cc/f6z to eaz your calculations

OpenStudy (agent0smith):

Ah, yes hyperphysics is handy that way.

OpenStudy (agent0smith):

@julliemyers first convert your times in minutes into seconds. Then use the formula for period (in seconds), using your length (in cm): \[T = 2 \pi \sqrt{\frac{ L }{ 980}}\]

OpenStudy (agent0smith):

There'll be several reasons why your experimental times are off from the theoretical times - friction, air resistance, human error (in timing etc). The formula for period is also simplified - it assumes the string has no mass, which of course isn't true in a real pendulum. If the mass of the bob is much heavier than the mass of the string, your experimental times will be more accurate.

OpenStudy (agent0smith):

Well, let me know if you need any more help/clarification.

OpenStudy (agent0smith):

Idk if you still need help on this, but your "Experimental Period" in the last table, should be the same as the "period of pendulum" from your other two tables. Experimental period just means what period you got from your experiment's results. Convert all the minutes into seconds. The "length of pendulum" is the length of the string you used, which looks like 25.6cm. Theoretical period is what you get from the formula, by using the length 25.6cm, and the period will be in seconds.

OpenStudy (anonymous):

thanks but i alredy turned it iin

OpenStudy (anonymous):

@julliemyers what was your answers for this or how did you get it

OpenStudy (anonymous):

@agent0smith do you know what it is or are you just looking

OpenStudy (agent0smith):

^if you read through it you'll see I answered it already.

OpenStudy (anonymous):

oh sorry lol